OpenStudy (undercover):

8x+y=-16 -3x+y=-5 solve the systems by elmination Name the solution as an ordered pair

OpenStudy (undercover):

it is only 1 question

OpenStudy (anonymous):

I don't know if this is right. but 8x+y=16 -8x=y=16-8x. the 8's cancel out. so it's Y=16-8x?

OpenStudy (anonymous):

I have a solution of 700ml of 30% acid, but need a solution of 50% acid. How many ml of a 100% acid solution should be added